# Opinion Dynamics for Utility Maximizing Agents: Exploring the Impact of Resource Penalty

### v1

@@ 7 Simulations
$\mathbf{w}$ $\approx\begin{bmatrix}13.42&14.30&12.84&8.38&7.82&16.32\end{bmatrix},$
@@
$\boldsymbol{\sigma}$ $\approx\begin{bmatrix}1.35&9.49&7.99&6.09&2&6.59\end{bmatrix}\times 10^{6}.$
@@
$\mathbf{w}$ $\approx\begin{bmatrix}2.76&16.73&2.77&11.76&7.32&16.14\end{bmatrix}$
$\mathbf{r}$ $\approx\begin{bmatrix}503.78&489.59&877.05&353.14&449.44&963.53\end{bmatrix}$
@@
$\mathbf{m}\approx\begin{bmatrix}-11.13&15.53&-10.47&6.12&3.36&6.92\end{bmatrix},$
which violates the necessary condition for consensus given in Theorem 5.1. From Figure 1(b), it can be seen that every agent’s opinion converges to a value in the set $\mathcal{M}\approx[-11.13,15.53]$ which verifies the results of Theorem 4.5 and Proposition 4.7.
Figure 2(a) demonstrates the oscillatory behavior exhibited by opinions of two agents under (22). The model parameters for this case are, $\mathbf{z_{0}}\approx\begin{bmatrix}-0.0349&-0.0039\end{bmatrix}$, $\mathbf{c}=\begin{bmatrix}-20&4.99\end{bmatrix}$, $\mathbf{w}=\begin{bmatrix}3&4\end{bmatrix}$, $\mathbf{p}=\begin{bmatrix}0&0\end{bmatrix}$ and $\mathbf{r}=\begin{bmatrix}10&5\end{bmatrix}$. Consider (22) with the above parameter values and let $c_{2}$ be the bifurcation parameter. Notice that since $p_{1}=p_{2}=0$, neutral consensus $\mathbf{z}^{*}=\mathbf{0}\in\mathcal{E}\>,\>\forall c_{2}\in\mathbb{R}$. These values satisfy the assumptions of Theorem 6.4 ensuring the existence a Hopf bifurcation at $c_{2}^{*}=5$. The eigenvalues of the Jacobian $\mathbf{J}(\mathbf{z}^{*},c_{2}^{*})$ are ±$\pm 0.0667j$. As a result, a family of periodic orbits bifurcates out of the neutral consensus equilibrium at $c_{2}^{*}$ and the periodic solution shown in Figure 2(a) is a member of this family corresponding to $c_{2}=4.99<c_{2}^{*}$. Figure 2(b) shows the corresponding trajectory in the phase plane intersecting the ellipse defined in (23). These simulations support the claim of Lemma 6.2 and its interpretation made in Remark 6.3.
@@ 8 Conclusions
We proposed a non-linear model of opinion dynamics to capture the effect of heterogeneous resources available to the agents on their opinions. In contrast to our prior work [26], we dealt with general social networks with (possibly) antagonistic relations. We showed ultimate boundedness of opinions and provided sufficient conditions for the dynamics to have a globally exponentially stable equilibrium point. We also provided necessary and sufficient condition for the existence of a consensus equilibrium and quantified social dominance at consensus. Further, we showed that the set of Nash equilibria of the opinion formation game is a subset of the set of equilibrium points of the dynamics and provided conditions for these two sets to coincide. In the absence of antagonistic relations, we gave stronger results. We quantified the quality of the Nash equilibria with respect to two commonly used prices of anarchy ($\mathsf{PoA}$), provided bounds on these $\mathsf{PoA}$’s in terms of the satisfaction ratios and proved that converging to a consensus equilibrium is a socially optimal outcome. Finally, we analyzed the periodic behavior of opinions exhibited by the proposed dynamics for the case of two agents. We provided necessary conditions for periodic solutions to exist and sufficient conditions for a Hopf bifurcation to occur at the consensus equilibrium.
Future research directions include extensions of the model to a multi-topic scenario, analysis of periodic behavior exhibited by opinions in the presence of antagonistic relations for a general $n$-agent case, and exploration of a more general class of utility functions and resource penalty functions.
@@
$[\mathbf{J}(\mathbf{z})]_{ij}=\begin{cases}\displaystyle\frac{-w_{i}r_{i}}{B}-\left(\sum_{k\in\mathcal{V}\setminus\left\{i\right\}}\frac{a_{ik}r_{k}}{B}\right)-\frac{3z_{i}^{2}}{r_{i}}\,,&\mathrm{if}\,i=j;\\ \displaystyle\frac{a_{ij}r_{j}}{B},&\mathrm{if}\,i\neq j.\end{cases}$
@@
$\mu_{\infty}(\mathbf{J}(\mathbf{z}))=\max_{i\in\mathcal{V}}\bigg{[}-\frac{w_{i}r_{i}}{B}+\sum_{k\in\mathcal{N}_{i}^{e}}\frac{2\,|a_{ik}|\,r_{k}}{B}-\frac{3z_{i}^{2}}{r_{i}}\bigg{]}.$
@@
Notice that under Assumption (A3), the dynamics (2) reduces to (4). First, we prove positive invariance of $\mathcal{M}^{n}$ under (4) by inspecting the vector field at $\mathbf{z}$ on the boundary of $\mathcal{M}^{n}$. Consider any $i\in\mathcal{V}$ such that $z_{i}=m_{\max}$. From (7), we see that $S_{i}(z_{i})\leq 0$ and $C_{i}(z_{i})\leq 0$, as $m_{i}\leq m_{\max}$ and $\bar{z}_{i}$ is a convex combination of $\{z_{j}\}_{j=1}^{n}$ and $\mathbf{z}\in\mathcal{M}^{n}$. Thus, $f_{i}(\mathbf{z})\leq 0$. Similarly, for any $i\in\mathcal{V}$ such that $z_{i}=m_{\min}$, $f_{i}(\mathbf{z})\geq 0$. This implies that $\mathcal{M}^{n}$ is positively invariant under (4).
Now, we show that $\mathbf{z}(t)$ converges to $\mathcal{M}^{n}$. Note that Assumption (A3) is a special case of Assumption (A4) and hence Theorem 4.5 guarantees that there is a unique globally exponentially stable equilibrium point $\mathbf{z}^{*}\in\mathcal{E}$. We know that under Assumption (A3), $\bar{z}_{i}$ is a convex combination of $\{z_{k}\}_{k\in\mathcal{N}_{i}}$ and that (7) holds. So, $\forall\mathbf{z}\notin\mathcal{M}^{n}$, $\exists\,i\in\mathcal{V}$ such that $\dot{z}_{i}\neq 0$. Hence, the unique globally exponentially stable equilibrium point $\mathbf{z}^{*}\in\mathcal{M}^{n}$ and as a result, $\mathbf{z}(t)$ converges to $\mathcal{M}^{n}$.
@@
A further consequence of (S1) is that for $i\in\mathcal{V}_{\max}$, if $\exists j\in\mathcal{N}_{i}\cap(\mathcal{V}\setminus\mathcal{V}_{\max})$ then $\bar{z}_{i}^{*}<m_{\max}$ and hence $z_{i}^{*}<m_{\max}$. We can use this fact along with (S1) and (S2) to also reason that for $i\in\mathcal{V}_{\max}$, $z_{i}^{*}<m_{\max}$ if and only if there is a directed walk in $\mathsf{G}$ from $j\in(\mathcal{V}\setminus\mathcal{V}_{\max})$ to $i$. We can make similar observations corresponding to $m_{\min}$. From all these observations, we can finally say that
@@
From the discussion just above Lemma 6.2, it is necessary that at least one of $c_{1}$ or $c_{2}<0$ for periodic solutions to exist. Further, it is also necessary that both conformity weights are non-zero. Otherwise, from Lemma 4.3, the socially closed agent’s opinion would converge to its own $m_{i}$ value. The opinion of the contrarian agent then would always converge to an equilibrium. Now, it remains to show that $(c_{2}r_{1}+w_{1}r_{1}+w_{2}r_{2}+c_{1}r_{2})<0$ is necessary for existence of periodic solutions. We prove the inverse using Bendixson’s criteria. Note the divergence of the vector field (22) at $\mathbf{z}\in\mathbb{R}^{2}$ is,
$\mathsf{div}(\mathbf{f}(\mathbf{z}))=-\frac{(w_{1}r_{1}+w_{2}r_{2}+c_{1}r_{2}+c_{2}r_{1})}{B}-\frac{3z_{1}^{2}}{r_{1}}-\frac{3z_{2}^{2}}{r_{2}}.$
Suppose $(c_{2}r_{1}+w_{1}r_{1}+w_{2}r_{2}+c_{1}r_{2})>0$. Then, $\mathsf{div}(\mathbf{f}(\mathbf{z}))<0$, $\forall\mathbf{z}\in\mathbb{R}^{2}$. Thus, by Bendixson’s criteria [30, Lemma 2.2] there are no periodic orbits lying entirely in $\mathbb{R}^{2}$.
@@
Under the assumption $m_{1}=m_{2}=m$, it follows from Theorem 5.1 that $\mathbf{z}^{*}=m\mathbf{1}\in\mathcal{E}\>;\>\forall c_{2}\in\mathbb{R}$. The Jacobian of (22) evaluated about $\mathbf{z}^{*}$ is,
$\mathbf{J}(\mathbf{z}^{*},c_{2})=\begin{bmatrix}(-\kappa_{1}-d)&d\\ gc_{2}&(-\kappa_{2}-gc_{2})\end{bmatrix}.$

### v2

@@ 7 Simulations
$\mathbf{w}$ $\approx\begin{bmatrix}1.09&2.98&2.59&1.82&1.01&2.65\end{bmatrix},$
@@
$\boldsymbol{\sigma}$ $\approx\begin{bmatrix}345.5&2424.4&2042.5&1555.1&510.4&1684.8\end{bmatrix}.$
@@
$\mathbf{w}$ $\approx\begin{bmatrix}0.22&3.48&0.56&2.56&0.95&2.62\end{bmatrix}$
$\mathbf{r}$ $\approx\begin{bmatrix}317.43&814.54&789.07&852.26&505.64&635.66\end{bmatrix}$
@@
$\mathbf{m}\approx\begin{bmatrix}-8.78&17.14&-10.10&6.56&3.38&6.81\end{bmatrix},$
which violates the necessary condition for consensus given in Theorem 5.1. From Figure 2(b), it can be seen that every agent’s opinion converges to a value in the set $\mathcal{M}\approx[-10.10,17.14]$ which verifies the results of Theorem 4.3 and Proposition 4.5.
Figure 3(a) demonstrates the oscillatory behavior exhibited by opinions of two agents under (17). The model parameters for this case are, $\mathbf{z_{0}}\approx\begin{bmatrix}-0.0349&-0.0039\end{bmatrix}$, $[a_{12}\>\>a_{21}]\approx\begin{bmatrix}-6.6667&3.3267\end{bmatrix}$, $\mathbf{w}\approx\begin{bmatrix}2&1.3333\end{bmatrix}$, $\mathbf{p}=\begin{bmatrix}0&0\end{bmatrix}$ and $\mathbf{r}=\begin{bmatrix}10&5\end{bmatrix}$. Consider (17) with the above parameter values and let $a_{21}$ be the bifurcation parameter. Notice that since $p_{1}=p_{2}=0$, neutral consensus $\mathbf{z}^{*}=\mathbf{0}\in\mathcal{E}\>,\>\forall a_{21}\in$. These values satisfy the assumptions of Theorem 6.4 ensuring the existence a Hopf bifurcation at $a_{21}^{*}=3.3333$. The eigenvalues of the Jacobian $\mathbf{J}(\mathbf{z}^{*},a_{21}^{*})$ are ±$\pm 0.6667j$. As a result, a family of periodic orbits bifurcates out of the neutral consensus equilibrium at $a_{21}^{*}$ and the periodic solution shown in Figure 3(a) is a member of this family corresponding to $a_{21}=3.3267<a_{21}^{*}$. Figure 3(b) shows the corresponding trajectory in the phase plane intersecting the ellipse defined in (18). These simulations support the claim of Lemma 6.2 and its interpretation made in Remark 6.3.
@@ 8 Conclusions
We proposed a non-linear model of opinion dynamics to capture the effect of heterogeneous resources available to the agents on their opinions. In contrast to our prior work [31], we dealt with general social networks with (possibly) antagonistic relations. We showed ultimate boundedness of opinions and provided sufficient conditions for the dynamics to have a globally exponentially stable equilibrium point. We also provided necessary and sufficient condition for the existence of a consensus equilibrium and quantified social dominance at consensus. Further, we showed that the set of Nash equilibria of the opinion formation game is a subset of the set of equilibrium points of the dynamics and provided conditions for these two sets to coincide. In the absence of antagonistic relations, we gave stronger results. We quantified the quality of the Nash equilibria with respect to two commonly used prices of anarchy ($\mathsf{PoA}$), provided bounds on these $\mathsf{PoA}$’s in terms of the satisfaction ratios and proved that converging to a consensus equilibrium is a socially optimal outcome. Finally, we analyzed the periodic behavior of opinions exhibited by the proposed dynamics for the case of two agents. We provided necessary conditions for periodic solutions to exist and sufficient conditions for a Hopf bifurcation to occur at the consensus equilibrium. Future research directions include extensions of the model to a multi-topic scenario, analysis of periodic behavior exhibited by opinions in the presence of antagonistic relations for a general $n$-agent case, and exploration of a more general class of utility functions and resource penalty functions.
@@
$[\mathbf{J}(\mathbf{z})]_{ij}=\begin{cases}\displaystyle-w_{i}-\left(\sum_{k\in\mathcal{V}\setminus\left\{i\right\}}a_{ik}\right)-\frac{3z_{i}^{2}}{r_{i}}\,,&\mathrm{if}\,i=j;\\ \displaystyle a_{ij},&\mathrm{if}\,i\neq j.\end{cases}$
@@
$\mu_{\infty}(\mathbf{J}(\mathbf{z}))=\max_{i\in\mathcal{V}}\bigg{[}-w_{i}+\sum_{k\in\mathcal{N}_{i}^{e}}2|a_{ik}|-\frac{3z_{i}^{2}}{r_{i}}\bigg{]}.$
@@
Notice that under Assumption (A3), the dynamics (2) reduces to (4). First, we prove positive invariance of $\mathcal{M}^{n}$ under (4) by inspecting the vector field at $\mathbf{z}$ on the boundary of $\mathcal{M}^{n}$. Consider any $i\in\mathcal{V}$ such that $z_{i}=m_{\max}$. From (7), we see that $S_{i}(z_{i})\leq 0$ and $C_{i}^{+}(\mathbf{z})\leq 0$, as $m_{i}\leq m_{\max}$ and $\bar{z}_{i}$ is a convex combination of $\{z_{j}\}_{j=1}^{n}$ and $\mathbf{z}\in\mathcal{M}^{n}$. Thus, $f_{i}(\mathbf{z})\leq 0$. Similarly, for any $i\in\mathcal{V}$ such that $z_{i}=m_{\min}$, $f_{i}(\mathbf{z})\geq 0$. This implies that $\mathcal{M}^{n}$ is positively invariant under (4).
Now, we show that $\mathbf{z}(t)$ converges to $\mathcal{M}^{n}$. Note that Assumption (A3) is a special case of Assumption (A4) and hence Theorem 4.3 guarantees that there is a unique globally exponentially stable equilibrium point $\mathbf{z}^{*}\in\mathcal{E}$. We know that under Assumption (A3), $\bar{z}_{i}$ is a convex combination of $\{z_{k}\}_{k\in\mathcal{N}_{i}}$ and that (7) holds. So, $\forall\mathbf{z}\notin\mathcal{M}^{n}$, $\exists\,i\in\mathcal{V}$ such that $\dot{z}_{i}\neq 0$. Thus, the unique globally exponentially stable equilibrium point $\mathbf{z}^{*}\in\mathcal{M}^{n}$ and hence $\mathbf{z}(t)$ converges to $\mathcal{M}^{n}$.
@@
A further consequence of (S1) is that for $i\in\mathcal{V}_{\max}$, if $\exists j\in\mathcal{N}_{i}\cap(\mathcal{V}\setminus\mathcal{V}_{\max})$ then $\bar{z}_{i}^{*}<m_{\max}$ and hence $z_{i}^{*}<m_{\max}$. This fact along with (S1) and (S2) imply that for $i\in\mathcal{V}_{\max}$, $z_{i}^{*}<m_{\max}$ if and only if there is a directed walk in $\mathsf{G}$ from $j\in(\mathcal{V}\setminus\mathcal{V}_{\max})$ to $i$. We can make similar observations corresponding to $m_{\min}$. From all these observations, we can finally say that
@@
From the discussion just above Lemma 6.2, it is necessary that at least one of $a_{12}$ or $a_{21}<0$ for periodic solutions to exist. Further, it is also necessary that both influence weights are non-zero. Otherwise, from Lemma 4.2, the socially closed agent’s opinion would converge to its own $m_{i}$ value. The opinion of other agent then would always converge to an equilibrium. Now, it remains to show that $(a_{21}+w_{1}+w_{2}+a_{12})<0$ is necessary for existence of periodic solutions. We prove the inverse using Bendixson’s criteria. Note the divergence of the vector field (17) at $\mathbf{z}\in^{2}$ is,
$\mathsf{div}(\mathbf{f}(\mathbf{z}))=-(w_{1}+w_{2}+a_{12}+a_{21})-\frac{3z_{1}^{2}}{r_{1}}-\frac{3z_{2}^{2}}{r_{2}}.$
Suppose $(a_{21}+w_{1}+w_{2}+a_{12})>0$. Then, $\mathsf{div}(\mathbf{f}(\mathbf{z}))<0$, $\forall\mathbf{z}\in^{2}$. Thus, by Bendixson’s criteria [37, Lemma 2.2] there are no periodic orbits lying entirely in 2.
@@
Under the assumption $m_{1}=m_{2}=m$, it follows from Theorem 5.1 that $\mathbf{z}^{*}=m\mathbf{1}\in\mathcal{E}\>;\>\forall a_{21}\in$. The Jacobian of (17) evaluated about $\mathbf{z}^{*}$ is,
$\mathbf{J}(\mathbf{z}^{*},a_{21})=\begin{bmatrix}(-\kappa_{1}-a_{12})&a_{12}\\ a_{21}&(-\kappa_{2}-a_{21})\end{bmatrix}.$